# Which Viewpoint Shows it Best? Language for Weakly Supervising View Selection in Multi-view Videos

### v1

@@ Abstract
Given a multi-view video, which viewpoint is most informative for a human observer? Existing methods rely on heuristics or expensive “best-view" supervision to answer this question, limiting their applicability. We propose a weakly supervised approach that leverages language accompanying an instructional multi-view video as a means to recover its most informative viewpoint(s). Our key hypothesis is that the more accurately an individual view can predict a view-agnostic text summary, the more informative it is. To put this into action, we propose a framework that uses the relative accuracy of view-dependent caption predictions as a proxy for best view pseudo-labels. Then, those pseudo-labels are used to train a view selector, together with an auxiliary camera pose predictor that enhances view-sensitivity. During inference, our model takes as input only a multi-view video—no language or camera poses—and returns the best viewpoint to watch at each timestep. On two challenging datasets comprised of diverse multi-camera setups and how-to activities, our model consistently outperforms state-of-the-art baselines, both with quantitative metrics and human evaluation.
@@ 1 Introduction
Figure 1: Main idea: given multi-view instructional videos, we aim to learn a view selection model that can identify the best view for learning how to perform the activity shown in the videos, in the absence of best view labels. To achieve this, we compare each estimated view-dependent caption to the view-agnostic ground-truth video narration of the human activity, and use their respective accuracies as a proxy for view quality. These quality scores then serve as pseudo-labels for learning to select the most informative view. In this example, the $i^{\text{th}}$ view most clearly shows all entities involved in the activity—the wheel and the person’s hands, and how they interact—and hence, produces a caption that best matches the ground-truth, making it a positive pseudo-label for view selection.
@@
Shooting video with multiple cameras provides a holistic view of the activity taking place in a scene, by capturing it from different locations and angles, and multi-view video is developing as a new frontier in computer vision research [85, 91, 65, 33, 44], especially in instructional settings [64, 33]. However, multi-view videos are generally not suitable for direct human consumption [21]: digesting multiple views at once imposes a high cognitive burden. Thus, in practice, the status quo is to orchestrate view selection in how-to videos manually with either active camera work or post-production video editing tools, which is time consuming and tedious.
@@
Conscious of these shortcomings, we propose to learn view selection in multi-view instructional videos in the absence of best view labels. Towards that goal, we hypothesize that view-agnostic natural language descriptions of the activity shown in the videos [33, 50, 44]—commonly referred to as “narrations"111 Narrations in multi-view datasets [33, 50, 44] are produced by human annotators who watch all views and write down a view-independent description of how the activity is performed, and in the wild they correspond to the “how-to” descriptions spoken by a person demonstrating a task [64]. can act as a source of weak supervision. Specifically, our core idea is that for any multi-view video clip, the viewpoint that is most predictive of such a narration is likely to be the most informative of the activity, and hence, can be pseudo-labeled as the best view for training a view selector. For example, given a multi-view video of a person repairing a bike (Figure 1), independent captions on each view will emphasize different visible components of the scene (the wheel, the person’s hands, other objects in the scene, etc.); the caption most aligned with the view-agnostic narration the person removes the rear wheel with both hands" indicates which view is most informative for the whole activity content in that clip. Unlike explicit best-view labels, the vision-language annotations that fuel today’s captioners are open-world, versatile, and widely available.
To validate our hypothesis, we design a novel framework composed of two key elements: a best view pseudo-labeler, and a best view selector. See Fig. 1. The pseudo-labeler automatically generates best view pseudo-labels for a multi-view video during training, by using off-the-shelf video captioners [109, 58] to score and rank views on the basis of how well the predicted narration from a view matches the view-agnostic ground-truth narration. The selector takes a multi-view video as input, and predicts the best view labels. During training, the selector also solves an auxiliary task of predicting the relative camera pose between different views, to increase its view-sensitivity and improve its selection accuracy. At inference, our model requires as input only a multi-view video, but no language or camera poses.
@@ 2 Related work
Temporal video summarization [71, 81, 75, 8, 68, 39] creates a synopsis of a long video by identifying and stitching together its most representative clips. Early unsupervised methods use hand-crafted features, optimizing for metrics like object saliency [62] and motion attention [70], while recent efforts [38, 31, 60, 57] use labeled data [86, 37] for supervised training. To mitigate the supervision cost, recent work leverage web priors [52, 53, 10] to target settings with insufficient or unreliable annotations [52, 102, 86, 73] or unpaired data [80, 105, 68]. Different from all of the above, we tackle this problem of label scarcity in the context of view selection in multi-view videos, a distinct task from video summarization. In multi-video summarization, the input videos are either captured with multi-view cameras in indoor surveillance [72, 28] or street [25, 110, 28] settings, or grouped using shared visual concepts [71, 63, 97, 82, 18]. Such methods can select multiple views at the same time, and hence, are not applicable in our setting. Unlike all video summarization methods, where the goal is to create a sparse temporal summary of one or more videos, our task requires choosing the appropriate camera view at each time step (clip).
@@ 3 Approach
Figure 2: (a) Our model uses language guidance to train a view-selector for multi-view instructional video, such that the chosen views help best understand the shown activity. To do so, we first generate best view pseudo-labels during training by leveraging clip narrations, where each narration is a view-agnostic and detailed description of the activity. Specifically, given a training clip, we use off-the-shelf video captioners to predict a narration estimate per view, score the views by comparing their narration estimates to the ground-truth, and finally rank the views to generate a best view pseudo-label for the clip. Given the multi-view clip, our view classifier (bottom-left) encodes it with a visual encoder, and predicts a pseudo-label estimate. We also solve an auxiliary task of relative camera pose prediction (bottom-right) that increases the view sensitivity of the classifier. (b) Examples of predicted narrations, and the ranks and scores of the views per our pseudo-labeler, shown alongside ground-truth view-agnostic narrations.
@@ 3.1 View selection task
Each clip thus has a ground-truth view-agnostic narration $\mathcal{N}_{m}^{*}$. We stress that each $\mathcal{N}_{m}^{*}$ captures the activity as viewed from any and all angles; it is view-independent. The descriptions include details about actions taken by the camera wearer as well as relevant events from the environment and important objects. We aim to use these narrations to train a model $\mathcal{F}$ that, given the video $V$, predicts a viewpoint sequence i.e., $\mathcal{F}(V)=\mathcal{B}^{*}$. We stress that the narrations are available only for training videos, not at test time.
@@ 3.2 Sourcing best view pseudo-labels for training
We devise our pseudo-labeler $L$ to generate the best-view pseudo-label for a training clip $\mathcal{V}_{m}$ by first predicting the narration for each view separately, and then scoring the views by comparing their predicted narrations to the view-agnostic ground-truth narration $\mathcal{N}_{m}$. [^2] For simplicity, we omit the clip index $m$ from subscript, henceforth. We aggregate results over multiple independent captioners in order to bolster robustness, essentially smoothing over their outputs. To this end, we use $K$ off-the-shelf video captioning models: 1) Video-Llama [109] with Llama2 [93] LLM decoder, 2) Video-Llama [109] with Vicuna [93] LLM decoder, and 3) VideoChat2 [58]. See Fig. 2(a) top.
@@
Essentially, our pseudo-labeler uses multiple strong captioners to rank views based on the accuracy of their predicted narrations, and achieves consensus on the top ranked views, thereby automatically producing high-quality best view pseudo-labels. See Fig. 2(b) for examples.
@@ 4 Experiments
We overview all setup details then provide results.
@@ 4.1 Experimental setup
We evaluate our model on two multi-view instructional video datasets: Ego-Exo4D [33] and LEMMA [50] [^3] Note that other instructional video datasets like HowTo100M [64], CrossTask [112], and COIN [92] are single-view and thus not suitable.. Ego-Exo4D contains both physical (e.g. basketball, dancing) and procedural (e.g. cooking, bike repair) activities with each video containing 5 time-synced views—one is egocentric (ego) and recorded by the participant with a headworn camera, and the remaining views are exocentric (exo) and recorded with stationary cameras kept around the scene. LEMMA has videos of people performing household activities (e.g. making juice, watering plant), where each video has the participant’s ego view and an exo view from a fixed camera placed facing the activity. Both have narrations: the Ego-Exo4D annotators provide temporally dense written descriptions of the participants’ actions and relevant objects involved in the activity, while the LEMMA annotators describe action verbs and objects being interacted with using a pre-defined vocabulary. These two datasets let us evaluate different exo view setups (multiple in EgoExo-4D vs. single in LEMMA) and diverse activity scenarios (physical and procedural in Ego-Exo4D vs. household in LEMMA). Ego-Exo4D and LEMMA provide a total of 86 and 20 hours of video data, resulting in 648665 and 63538 clip-narration pairs, respectively.
@@
We use $K=3$ captioners in our pseudo-labeler: Video-Llama [109] with Llama-2-Chat [93] or Vicuna [16] as the LLM decoder, and VideoChat2 [58], respectively, and the CIDEr [95] captioning metric to score views. We finetune the captioners on our datasets before using them to score views. On the basis of a disjoint validation set, we set the bin size to $30$ degrees for generating relative camera pose labels (cf. Sec. 3.3), and the weight on the pose prediction loss to $w=0.5$.
@@
  • Longest-caption: a baseline that uses our finetuned Video-Llama captioner [109] to predict narrations for each input view, and selects the one that produces the longest narration as the best view. Intuitively, here narration length is used a proxy for informativeness. On the contrary, our model does not infer narration on test data.
While the first set of baselines are naive heuristics to test if intelligent view selection is even necessary, the second set accounts for the prior that visibility of people and person-object interactions are strongly linked to the informativeness of a view, and the third represents the most relevant existing methods in the literature. Finally, the longest-narration baseline offers an alternative, more naive way to incorporate language for our task.
@@ 4.2 Results
Our model significantly outperforms all baselines across metrics, on both datasets. It shows our idea of pseudo-labeling by leveraging the view-dependent quality of predicted narrations is effective in practice. Furthermore, our model’s improvement over the heuristics illustrates that our language-guided training facilitates complex but essential reasoning about the interplay between human actors and interacting objects, more than what is possible with hand and object, or body pose detectors. Finally, our superior performance on both datasets underlines the efficacy of our design and its ability to generalize to different activity types—both physical and non-physical in Ego-Exo4D vs. household in LEMMA— and camera setups—single vs. multiple exo cameras in LEMMA and Ego-Exo4D, respectively.
@@
In our view selection assessment, we find that our selected views are preferred significantly more than the two top baselines’ on both datasets. We achieve a strong win rate boost of at least $\sim$11over the baselines across different evaluation scenarios. The only exception is our model vs. Hand-object [15] on LEMMA, where the improvement margin is a more modest $\sim$2%, likely because the view that best shows hand-object interactions is a reasonably good view for showing LEMMA-style activity (cooking, watering plant, etc.). These results show that our weakly-supervised method is better capable of implicitly modeling human view preference and using this model to perform more accurate view selection. Finally, our consistent performance gains on human evaluation reinforce our automatic evaluation metrics, and all model analysis we do using them.
@@
Figure 3: Left: sample successful predictions by our view selector. For each clip, our model chooses the view that shows the action, and the objects and body parts involved in it, most clearly, and hence, is most suitable for activity understanding. Right: Sample failure cases for our model, where there are multiple high-quality views that differ only in certain nuances, which are discernible by a human but not our model trained through narration guidance. Whereas humans prefer a view that better captures the direction of the ball towards the camera-wearer in sample 1, or shows the full backward motion of the dancers in sample 2, our model learns to choose a view that clearly shows all humans, objects and actions mentioned in the narration. In such cases, our model predictions might not match human preferences.
Fig. 3 (left) shows our model’s success cases. Note how our model chooses views that clearly show all the essential elements of the actions, including the different entities involved and and their motion. E.g., in clip 2, our model chooses the ego view as it clearly shows the horizontal knife motion and vegetables on the cutting board, whereas in clip 3, it selects an exo view that best shows the joined hands and the sideways movement of the dancers. Thus, depending on the activity, our model adaptively chooses a view that accurately shows its important details.
We also observe two common failure types. The first type occurs when there are multiple high-quality views that differ only in certain intricate details, which are discernible by a human but not by our model due to its narration-guided training. In such cases, our model prediction might not match human preferences. See Fig. 3 (right). In the second one, our model chooses different views for very similar activities, which involve very similar types, positions and motion of relevant objects and body parts, even when the views are not equally good. This occurs possibly because our selector occasionally picks on spurious cues that are not as viewpoint-dependent as the activity itself but also do not help with its understanding.
@@
In Table 3 (top) we report our model ablation results on the large-scale Ego-Exo4D [33] dataset. Not predicting inter-view camera pose in training significantly hurts performance on all metrics except METEOR [6]. This shows that our pose predictor enhances the selector’s view sensitivity. Training our model with a standard cross-entropy loss by randomly a sample from our pseudo-label set, instead of our proposed loss (c.f. Sec. 3.4) hurts performance on N- and NC-IoU. This occurs possibly because the captioners in our pseudo-labeler sometimes hallucinate and add less informative views to the pseudo-label set, which when sampled causes training instability. Removing the rank aggregator to obtain pseudo-label consensus also negatively impacts performance on N- and NC-IoU, as having multiple captioners vote on the best view reduces captioning noise and improves pseudo-label quality. Finally, removing all three components consistently degrades performance across metrics, showing that the model needs at least one component to perform well on different metrics.

### v2

@@ Abstract
Given a multi-view video, which viewpoint is most informative for a human observer? Existing methods rely on heuristics or expensive “best-view" supervision to answer this question, limiting their applicability. We propose a weakly supervised approach that leverages language accompanying an instructional multi-view video as a means to recover its most informative viewpoint(s). Our key hypothesis is that the more accurately an individual view can predict a view-agnostic text summary, the more informative it is. To put this into action, we propose a framework that uses the relative accuracy of view-dependent caption predictions as a proxy for best view pseudo-labels. Then, those pseudo-labels are used to train a view selector, together with an auxiliary camera pose predictor that enhances view-sensitivity. During inference, our model takes as input only a multi-view video—no language or camera poses—and returns the best viewpoint to watch at each timestep. On two challenging datasets comprised of diverse multi-camera setups and how-to activities, our model consistently outperforms state-of-the-art baselines, both with quantitative metrics and human evaluation. Project: http://vision.cs.utexas.edu/projects/which-view-shows-it-best.
@@ 1 Introduction
Figure 1: Main idea: given multi-view instructional videos, we aim to learn a view selection model that can identify the best view for seeing how to perform the activity shown in the videos, in the absence of best view labels. To achieve this, we compare each estimated view-dependent caption to the view-agnostic ground-truth video narration of the human activity, and use their respective accuracies as a proxy for view quality. These quality scores then serve as pseudo-labels for learning to select the most informative view. In this example, the 1st view most clearly shows all entities involved in the activity—the wheel and the person’s hands, and how they interact—and hence, produces a caption that best matches the ground-truth, making it a positive pseudo-label for view selection.
@@
Shooting a video with multiple cameras provides a holistic view of the activity taking place in a scene, by capturing it from different locations and angles, and multi-view video is developing as a new frontier in computer vision research [95, 101, 74, 37, 49], especially in instructional settings [73, 37]. However, multi-view videos are generally not suitable for direct human consumption [24]: digesting multiple views at once imposes a high cognitive burden. Thus, in practice, the status quo is to orchestrate view selection in how-to videos manually with either active camera work or post-production video editing tools, which is time consuming and tedious.
@@
Conscious of these shortcomings, we propose to learn view selection in multi-view instructional videos in the absence of best view labels. Towards that goal, we hypothesize that view-agnostic natural language descriptions of the activity shown in the videos [37, 56, 49]—commonly referred to as “narrations"111 Narrations in multi-view datasets [37, 56, 49] are produced by human annotators who watch all views and write down a view-independent description of how the activity is performed, and in the wild they correspond to the “how-to” descriptions spoken by a person demonstrating a task [73]., can act as a source of weak supervision. Specifically, our core idea is that for any multi-view video clip, the viewpoint that is most predictive of such a narration is likely to be the most informative of the activity, and hence, can be pseudo-labeled as the best view for training a view selector. For example, given a multi-view video of a person repairing a bike (Figure 1), independent captions on each view will emphasize different visible components of the scene (the wheel, the person’s hands, other objects in the scene, etc.); the caption most aligned with the view-agnostic narration the person removes the rear wheel with both hands" indicates which view is most informative for the whole activity content in that clip. Unlike explicit best-view labels, the vision-language annotations that fuel today’s captioners are open-world, versatile, and widely available.
To validate our hypothesis, we design a novel framework composed of two key elements: a best view pseudo-labeler, and a best view selector. The pseudo-labeler automatically generates best view pseudo-labels for a multi-view video during training, by using off-the-shelf video captioners [121, 64] to score and rank views on the basis of how well the predicted narration from a view matches the view-agnostic ground-truth narration. The selector takes a multi-view video as input, and predicts the best view labels. During training, the selector also solves an auxiliary task of predicting the relative camera pose between different views, to increase its view-sensitivity and improve its selection accuracy. At inference, our model requires as input only a multi-view video, but no language or camera poses.
@@ 2 Related work
Temporal video summarization [80, 91, 84, 9, 77, 43] creates a synopsis of a long video by identifying and stitching together its most representative clips. Early unsupervised methods use hand-crafted features, optimizing for metrics like object saliency [71] and motion attention [79], while recent efforts [42, 35, 66, 63] use labeled data [96, 41] for supervised training. To mitigate the supervision cost, recent work leverage web priors [58, 59, 11] to target settings with insufficient or unreliable annotations [58, 114, 96, 82] or unpaired data [90, 117, 77]. Different from all of the above, we tackle the problem of label scarcity in the context of view selection in multi-view videos, a distinct task from video summarization. In multi-video summarization, the input videos are either captured with multi-view cameras in indoor surveillance [81, 32] or street [29, 122, 32] settings, or grouped using shared visual concepts [80, 72, 109, 92, 19]. Such methods can select multiple views at the same time, and hence, are not applicable in our setting. Unlike all video summarization methods, where the goal is to create a sparse temporal summary of one or more videos, our task requires choosing the appropriate camera view at each time step (clip).
@@ 3 Approach
Figure 2: (a) Our model uses language guidance to train a view-selector for multi-view instructional videos, such that the chosen views help best understand the shown activity. To do so, we first generate best view pseudo-labels during training by leveraging clip narrations, where each narration is a view-agnostic and detailed description of the activity. Specifically, given a training clip, we use off-the-shelf video captioners to predict a caption per view, score the views by comparing their captions to the ground-truth narration, and finally rank the views to generate a best view pseudo-label for the clip. Given the multi-view clip, our view classifier (bottom-left) encodes it with a visual encoder, and predicts a pseudo-label estimate. We also solve an auxiliary task of relative camera pose prediction (bottom-right) that increases the view sensitivity of the classifier. (b) Examples of predicted narrations, and the ranks and scores of the views per our pseudo-labeler, shown alongside ground-truth view-agnostic narrations. Note that at inference time, there is no ground truth narration, just the video input.
@@ 3.1 View selection task
Each clip thus has a ground-truth view-agnostic narration $\mathcal{N}_{m}^{*}$. We stress that each $\mathcal{N}_{m}^{*}$ captures the activity as viewed from any and all angles; it is view-independent. The descriptions include details about actions taken by the camera wearer to the activity, and as well as relevant events from the environment and important objects. We aim to use these narrations to train a model $\mathcal{F}$ that, given the video $V$, predicts a viewpoint sequence i.e., $\mathcal{F}(V)=\mathcal{B}^{*}$. We stress that the narrations are available only for training videos, not at test time.
@@ 3.2 Sourcing best view pseudo-labels for training
We devise our pseudo-labeler $L$ to generate the best-view pseudo-label for a training clip $\mathcal{V}_{m}$ by first predicting the narration for each view separately, and then scoring the views by comparing their predicted narrations to the view-agnostic ground-truth narration $\mathcal{N}_{m}^{*}$. [^2] For simplicity, we omit the clip index $m$ from subscript, henceforth. We aggregate results over multiple independent captioners in order to bolster robustness, essentially smoothing over their outputs. To this end, we use $K$ off-the-shelf video captioning models: 1) Video-Llama [121] with Llama2 [103] LLM decoder, 2) Video-Llama [121] with Vicuna [103] LLM decoder, and 3) VideoChat2 [64]. See Fig. 2(a) top.
@@
Essentially, our pseudo-labeler uses multiple strong captioners to rank views based on the accuracy of their predicted narrations, and achieves consensus on the top ranked views, thereby automatically producing high-quality best view pseudo-labels. See Fig. 2(b) and Supp. for examples.
@@ 4 Experiments
We overview all setup details and then provide results.
@@ 4.1 Experimental setup
We evaluate our model on two multi-view instructional video datasets: Ego-Exo4D [37] and LEMMA [56] [^3] Note that other instructional video datasets like HowTo100M [73], CrossTask [124], and COIN [102] are single-view and thus not suitable.. Ego-Exo4D contains both physical (e.g. basketball, dancing) and procedural (e.g. cooking, bike repair) activities with each video containing 5 time-synced views—one is egocentric (ego) and recorded by the participant with a headworn camera, and the remaining views are exocentric (exo) and recorded with stationary cameras kept around the scene. LEMMA has videos of people performing household activities (e.g. making juice, watering plant), where each video has the participant’s ego view and an exo view from a fixed camera placed facing the activity. Both have narrations: the Ego-Exo4D annotators provide temporally dense written descriptions of the participants’ actions and relevant objects involved in the activity, while the LEMMA annotators describe action verbs and objects being interacted with using a pre-defined vocabulary. These two datasets let us evaluate different exo view setups (multiple in EgoExo-4D vs. single in LEMMA) and diverse activity scenarios (physical and procedural in Ego-Exo4D vs. household in LEMMA). Ego-Exo4D and LEMMA provide a total of 86 and 20 hours of video data, resulting in 648,665 and 63,538 clip-narration pairs, respectively. See Supp. for details.
@@
We use $K=3$ captioners in our pseudo-labeler: Video-Llama [121] with Llama-2-Chat [103] or Vicuna [17] as the LLM decoder, and VideoChat2 [64], respectively, and the CIDEr [107] captioning metric to score views. We finetune the captioners on our datasets before using them to score views. On the basis of a disjoint validation set, we set the bin size to $30$ degrees for generating relative camera pose labels (cf. Sec. 3.3), and the weight on the pose prediction loss to $w=0.5$. See Supp. for more details.
@@
  • Longest-caption: a baseline that uses our finetuned Video-Llama captioner [121] to predict captions for each input view, and selects the one that produces the longest narration as the best view. Intuitively, here caption length is used a proxy for informativeness. Recall that our model does not infer captions on test data.
While the first set of baselines are naive heuristics to test if intelligent view selection is even necessary, the second set accounts for the prior that visibility of people and person-object interactions are strongly linked to the informativeness of a view, and the third represents the most relevant existing methods in the literature. Finally, the Longest-caption baseline offers an alternative, more naive way to incorporate language for our task.
@@ 4.2 Results
Our model significantly outperforms all baselines across metrics, on both datasets. It shows our idea of pseudo-labeling by leveraging the view-dependent quality of predicted narrations is effective in practice. Furthermore, our model’s improvement over the heuristics illustrates that our language-guided training facilitates complex but essential reasoning about the interplay between human actors and interacting objects, more than what is possible with hand and object, or body pose detectors. Finally, our superior performance on both datasets underlines the efficacy of our design and its ability to generalize to different activity types—both physical and non-physical in Ego-Exo4D vs. household in LEMMA— and camera setups—single vs. multiple exo cameras in LEMMA and Ego-Exo4D, respectively. See Supp. for results on the single exo camera variant for Ego-Exo4D.
@@
In our view selection assessment, we find that our selected views are preferred significantly more than the two top baselines on both datasets. We achieve a strong win rate boost of at least $\sim$11over the baselines across different evaluation scenarios. The only exception is our model vs. Hand-object [16] on LEMMA, where the improvement margin is a more modest $\sim$2%, likely because the view that best shows hand-object interactions is a reasonably good view for showing LEMMA-style activity (cooking, watering plant, etc.). These results show that our weakly-supervised method is better capable of implicitly modeling human view preference and using this model to perform more accurate view selection. Finally, our consistent performance gains on human evaluation reinforce our automatic evaluation metrics, and all model analysis we do using them.
@@
Figure 3: Left: sample successful predictions by our view selector. For each clip, our model chooses the view that shows the action, and the objects and body parts involved in it, most clearly, and hence, is most informative. Right: Sample failure cases for our model, where there are multiple high-quality views that differ only in certain nuances, which are discernible by a human but not our model trained through narration guidance. Whereas humans prefer a view that better captures the direction of the ball towards the camera-wearer in sample 1, or shows the full backward motion of the dancers in sample 2, our model choose a view that shows all entities mentioned in the narration.
Fig. 3 (left) shows our model’s success cases. Note how our model chooses views that clearly show all the essential elements of the actions, including the different entities involved and and their motion. E.g., in clip 2, our model chooses the ego view as it clearly shows the horizontal knife motion and vegetables on the cutting board, whereas in clip 3, it selects an exo view that best shows the joined hands and the sideways movement of the dancers. Thus, depending on the activity, our model adaptively chooses a view that accurately shows its important details. See Supp. video for more examples.
We also observe two common failure types. The first type occurs when there are multiple high-quality views. See Fig. 3 (right). In the second one, our model chooses different views for very similar activities, which involve very similar types, positions and motion of relevant objects and body parts, even when the views are not equally good. This occurs possibly because our selector occasionally picks on spurious cues that are not as viewpoint-dependent as the activity itself but also do not help with its understanding.
@@
In Table 3 (top) we report our model ablation results on the large-scale Ego-Exo4D [37] dataset. Not predicting inter-view camera pose in training significantly hurts performance on all metrics except METEOR [7]. This shows that our pose predictor enhances the selector’s view sensitivity. Training our model with a standard cross-entropy loss by randomly choosing a sample from our pseudo-label set, instead of our proposed loss (c.f. Sec. 3.4) hurts performance on N- and NC-IoU. This occurs possibly because the captioners in our pseudo-labeler sometimes hallucinate and add less informative views to the pseudo-label set, which when sampled causes training instability. Removing the rank aggregator to obtain pseudo-label consensus also negatively impacts performance on N- and NC-IoU, as having multiple captioners vote on the best view reduces captioning noise and improves pseudo-label quality. Finally, removing all three components consistently degrades performance across metrics, showing that the model needs at least one component to perform well on different metrics.
See Supp. for additional ablations, and analyses of the view dependence of our visual features, our pseudo-labeler, and the impact of the rank of our selector’s sampled view on view selection performance.